You're betting on outcome of 2 dice–the complexity comes from how many different ways you can bet.
The goal is to predict what the dice will land on — either on a single roll or across a series of rolls. The more you understand the bets, the more tools you have.
11
2/36 = 5.56% · 17 to 1 against
| Result | Ways | Probability | Odds against |
|---|---|---|---|
| 2 | 1 | 1/36 = 2.78% | 35 to 1 |
| 3 | 2 | 2/36 = 5.56% | 17 to 1 |
| 4 | 3 | 3/36 = 8.33% | 11 to 1 |
| 5 | 4 | 4/36 = 11.11% | 8 to 1 |
| 6 | 5 | 5/36 = 13.89% | 6.2 to 1 |
| 7 | 6 | 6/36 = 16.67% | 5 to 1 |
| 8 | 5 | 5/36 = 13.89% | 6.2 to 1 |
| 9 | 4 | 4/36 = 11.11% | 8 to 1 |
| 10 | 3 | 3/36 = 8.33% | 11 to 1 |
| 11 | 2 | 2/36 = 5.56% | 17 to 1 |
| 12 | 1 | 1/36 = 2.78% | 35 to 1 |